OpenStudy (anonymous):

Simplify the expression (-2i)(8i)

OpenStudy (loser66):

(-2)*8 =?

OpenStudy (anonymous):

-16

OpenStudy (loser66):

i*i =?

OpenStudy (anonymous):

But what do I do with the i's?

OpenStudy (anonymous):

i^2

OpenStudy (loser66):

and i^2 = -1, right?

OpenStudy (loser66):

so (-16)*(-1) =?

OpenStudy (anonymous):

Oh I didn't know that. Okay so it's 16. Thank you!

OpenStudy (loser66):

np